A particle of length l having tension T and radius r vibrates with fundamental frequency f another wire of the same metal with length 2l and tension 2t and radius 2r will vibrate with funamental frequency
1-f
2-2f
3-f/2rt2
4-frt2/2

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago

m2=8m1
T=2t
l=2l
f new= sqrt(2*2/8)/2 f= f/2 sqrt(2)